Stuck and needs help with these math problems please help
1. Solve: x/3 = 3/x-8
2. Solve: 3x + 2/x + 2 - 1 = 2x
3. Solve: (x-2)(x-4)(x+1)<0
4. Solve: 3x/x-5 <5

Answer by stanbon(75887)

1. Solve: x/3 = 3/x-8
x(x-8) = 9
x^2-8x-9 = 0
(x-9)(x+1) = 0
x = 9 or x = -1
=========================
2. Solve: 3x + 2/x + 2 - 1 = 2x
(3x+2)/(x+2) - 1 = 2x
-----
3x+2 - (x+2) = 2x(x+2)
---
3x+2 -x-2 = 2x^2+4x
2x^2 +2x + 2 = 0
x^2 + x + 1 = 0
Use the Quadratic formula.
===============================
3. Solve: (x-2)(x-4)(x+1)<0
Draw a Number Line.
Plot: x = -1, x = 2, x = 4
---
Test a value in each interval.
Test x = -10: -*-*- < 0 true
Test x = 0 : -*-*+ < 0 false
Test x = 3 : +*-*+ < 0 true
Test x = 10 : +*+*+ < 0 false
----
Solution:(-oo,-1)U(2,4)
===========
Cheers,
Stan H.
===============
4. Solve: 3x/x-5 <5